### v1

@@ Abstract
Large Language Models (LLMs) and Large Multi-Modal Models (LMMs) have emerged as transformative tools in scientific research, yet their reliability and specific contributions to biomedical applications remain insufficiently characterized. In this study, we present ARtificial Intelligence research assistant for Expert-involved Learning (ARIEL), a multimodal dataset designed to benchmark and enhance two critical capabilities of LLMs and LMMs in biomedical research: summarizing extensive scientific texts and interpreting complex biomedical figures. To facilitate rigorous assessment, we create two open-source sets comprising biomedical articles and figures with designed questions. We systematically benchmark both open- and closed-source foundation models, incorporating expert-driven human evaluations conducted by doctoral-level experts. Furthermore, we improve model performance through targeted prompt engineering and fine-tuning strategies for summarizing research papers, and apply test-time computational scaling to enhance the reasoning capabilities of LMMs, achieving superior accuracy compared to human-expert corrections. We also explore the potential of using LMM Agents to generate scientific hypotheses from diverse multimodal inputs. Overall, our results delineate clear strengths and highlight significant limitations of current foundation models, providing actionable insights and guiding future advancements in deploying large-scale language and multi-modal models within biomedical research.
@@ 1 Introduction
The success of Foundation Models (FMs) [1], including Large Language Models (LLMs) and Large Multi-Modal Models (LMMs), has been demonstrated across various applications, garnering widespread public interest in research focused on these models [2, 3, 4, 5]. A very important target of an FM is to understand various modalities and model them to enable multimodal reasoning [6, 7] as humans. These human skills give us the ability to understand what is being conveyed in a multimodal context, while at the same time being able to externalize it with our linguistic system for processing and refining the content of the text [8]. In this work, we investigate how we can utilize FMs to help humans summarize long research papers as well as understand multimodal scientific context.
@@
There are various long documents in the biomedical domain, including examination reports, patient case data, research manuscripts, and others [9]. These documents may also include figures as extra information to help us understand the corresponding content. However, due to the explosion of biomedical data nowadays, physicians and researchers are struggling more and more to understand and process such documents in biomedical areas [10]. Also, the cost of hiring additional human resources to help process the data is high and therefore not optimal for the biomedical field [11, 12]. At the same time, there has been evidence [13, 14, 15] to show that humans tend to engage in complicated reasoning processes and expand their knowledge when answering realistic questions. For example, it is common to use multiple examination results to determine the condition [16] during the diagnosis from a physician. Therefore, how to measure and optimize the reasoning capacity of FMs for complex multimodal data is an important and practical research problem [17]. If an FM shows an excellent ability to summarize long texts as well as complicated figures with different panels, it will become a powerful research assistant in this field, which can shorten the time for physicians to make correct diagnoses and for researchers to understand the work of their peers.
Previous research has investigated the ability of LLMs to summarize non-medical documents including news [18], codes [19], and reviewer comments [20, 21]. Research has also been conducted to evaluate LLMs for summarizing short medical documents including clinical texts [12] and medical evidence [22]. Meanwhile, the ability of LMMs to summarize scientific figures has also been explored [23, 24]. These papers argue that some advanced models such as GPT-4 can outperform human experts in highlighting the abstract of certain documents or figures, while all models have their limitations due to either hallucinations or out-of-domain queries. Furthermore, most of the recent LLM or LMM evaluations have focused on comparisons among models but ignored the role of humans in evaluations. However, humans are very important references in evaluating the utilities of FMs for their purpose, and understanding the differences between FMs and humans can further ensure the reliability of these FMs for scientific research.
@@
As far as we know, comprehensive human-aware evaluations of LMMs for understanding long biomedical documents and biomedical figures remain unexplored. Although there are several LMM benchmarks on low-level knowledge quizzes, they cannot be applied to our problem due to the complexity of biomedical resources [25, 26, 27]. In addition, due to the rapid development of LMMs, we expect more up-to-date evaluations and benchmarks in this time-sensitive domain [28]. Most of the existing baselines also do not include human participators as an important metric. Researchers need to get fair and unbiased results carefully to avoid leakage of the test set. Furthermore, inspired by previous work focusing on generating scientific hypotheses from text-level information [29, 30], we are interested in understanding how multimodal biomedical data can help generate new research questions to accelerate scientific progress. Therefore, our LMM evaluations and improvement are challenging and meaningful, which can fill a gap in this advanced area.
@@
Here we present a novel dataset accompanied with analytical framework, named as ARIEL, for evaluating the understanding ability of LLMs and LMMs in the biomedical domain. Here, we focus on two basic tasks, i.e., long manuscript summarization (mainly related to COVID-19 and healthcare) and complicated figure comprehension (mainly related to multiomics) and provide several strategies for performance enhancement. To make a fair and comprehensive comparison, we manually constructed an open-source biomedical article summarization dataset as well as a biomedical figure understanding dataset. Our benchmark based on ARIEL not only systematically evaluates the performance of representative open- and closed-source LLMs and LMMs on both two tasks, but also offers human-aware evaluations from experts in the biomedical field. By analyzing the outputs and experts’ answers, we summarize the strengths and weaknesses of each model type, which enables us to explore how to leverage FMs to facilitate scientific inquires. We also test the performances of ARIEL in generating biomedical hypothesis with different frameworks. Overall, we believe that our multi-functional framework can significantly advance the applications of LLMs and LMMs in the healthcare domain.
@@ 2 Results
Overview of ARIEL To systemically assess the abilities of FMs in understanding research papers, we first constructed two separate datasets for both research manuscripts and research figures. These data were extracted from PubMed [31] and top-tier journals published recently and were not used for model training. We then set up a general pipeline for generating and automatically evaluating the outputs of both closed- and open-source FMs using these datasets. For the text summarization task, a language model receives a manuscript and a prompt as input. For the figure understanding task, a multimodal model takes as input the figure, its description, and a question prompt. In addition to automatic evaluation, we conducted a human evaluation (Human-Eval) with 12 doctoral-level biomedical researchers who judged model outputs and their sources. We then analyzed all experimental results to identify key gaps between current FM capabilities and the requirements of an effective AI research assistant. Finally, we propose initial strategies for using FMs to aid in summarizing documents, interpreting scientific figures, and generating scientific hypotheses. A summary of the ARIEL pipeline is shown in Figure 1.
@@
Figure 1: Landscape of ARIEL. (a) We designed a framework supporting 1. the evaluation of both open-source and closed-source LLMs and LMMs; and 2. the utilization of LLMs and LMMs for helping human researchers focus on both the text summarization task and the figure understanding task with new datasets. We also introduced numerical evaluation metrics and human-aware evaluation metrics for both tasks. The solid arrows in the panel indicate the steps that must be executed, and the dashed arrows indicate the steps that are optional. (b) Number of papers with top 10 topics in datasets used for biomedical document summarization. (c) Proportion of questions in different categories for scientific figure understanding.
Introduction of datasets. All of the questions collected in this dataset are open-ended. For the long document summarization task, we collected over 2000 manuscripts by using the API from PubMed. Compared with other datasets for document summarization evaluation [32, 33, 34], ARIEL contains large-scale information on biomedical domain specifically and scientific papers are more rigorous than other types of natural language for summarization. We also propose open-ended questions to make it more challenging than traditional QA format. We split them into training and testing datasets (To fine-tune a model, the validation dataset will be generated automatically from the training dataset by [35]). We visualize the topic statistics for all of the manuscripts in Figure 1 (b), which shows that COVID-19 and SARS-COV-2 are the two top topics for these research papers, followed by other topics such as healthcare and treatment. Therefore, our constructed datasets are centralized in the biomedical area and mainly focus on describing research related to epidemics of the respiratory system.
For the scientific figure understanding task, we carefully constructed ten figures with paired four to ten questions for each figure with 88 questions in total. Compared with other benchmarking datasets for testing scientific figure understanding [24, 25], our figures are from high-quality and peer-reviewed publications [23, 36, 37] to ensure a high signal-to-noise ratio, and we also provide category information for each open-ended question. These figures focus on describing the scientific discovery, model design, and benchmarking results for research in multi-omics (genomics, proteomics), which also overlap with our selected manuscripts in the first task. We also designed several important questions for the paired figure, and the questions covered aspects including figure summarization, color identification, model design, rank identification, recommendation, and insight discovery. The proportion of topics is shown in Figure 1 (c). These abilities need participators to perform reasoning analysis and are highly related to the main idea of scientific figures.
Regarding the experiments based on humans (Human-Eval), to reduce labor and collect high-quality feedback, we sampled ten papers from the testing dataset randomly in the first task and included all figures in the second task. We then assigned the questions and manuscripts (figures) to the human evaluators and collected their answers for analysis.
@@
Figure 2: Evaluations based on the text summarization task. (a) Pipeline of evaluation framework for this task. (b) Box plots of model performances by different evaluation metrics. (c) Averaged model performances by different evaluation metrics. Scores lower than 1.0 are annotated.
@@
Large-scale evaluations of long document summarization identify fine-tuned models as better performers. In this section, we focus on analyzing our evaluation results for long document summarization based on computational metrics. We considered both open-source LLMs (ChatGLM4 (9B) [38], Gemma1 (8B) [39], Llama 3.1 (8B) [40], Mistral (7B) [41]) and closed-source LLMs (GPT-3.5 [42], GPT-4 [43], Gemini 1.5 (Flash) [44], and Claude 3.5 (Sonnet)[45]) to assess the performances of models with different scales or training strategies. We selected the best model and fine-tuned it with Low-Rank Adaptation (LoRA) [46] to test if including more task-specific information can help the summarization ability of LLMs or not. To comprehensively compare the similarity between model-generated information and ground truth, we referred previous research [12, 47] and chose seven metrics, including BLEU, ROUGE-1, ROUGE-2, ROUGE-L, BERT score (BERT), F1-Radgraph, and MEDCON [48, 49, 50, 51, 47]. All of the metrics range from 0 to 100 and a higher value means a better output. Details of selected models and metrics are included in the Methods section, and the pipeline of our evaluation can be found in Figure 2 (a).
Based on Figure 2 (b), we compare the distribution and the variance of scores for all the metrics. We found that open-source models for the zero-shot inference mode generally performed worse than closed-source models under every metric, as we could not find high-score outputs from Gemma1, Llama 3.1, and Mistral. Moreover, ChatGLM4, Claude 3.5 (Sonnet), and GPT-4 are the best raw models in generating the summarization of long documents, which is consistent with previous investigations on using LLMs as reviewers [52]. Since ChatGLM4 performed better than other open-source models under every metric and had similar performance compared with closed-source LLMs, we tried different prompts to query it or fine-tuned it (as ChatGLM_lora), and evaluated these different modes. As shown in Extended Data Figure 2, different prompt engineering methods including meta prompting [53] and chain-of-thought (COT) [6] did not affect the model performances. Moreover, the fine-tuned version outperformed all closed-source LLMs, suggesting that the capacity of LLMs in summarizing the manuscript can be improved by training with domain-specific data, and small-scale models after fine-tuning can also become good task-specific performers. In Figure 2 (c), we display the average score of each metric for all the models, and the results are consistent with conclusions from Figure 2 (b).
@@
Figure 3: Comparisons between human participators and LLMs for the text summarization task. (a) Boxplots of average scores across all metrics by different participators. (b) Boxplots of participators’ performances by different evaluation metrics. (c) Case study based on the original abstract, human-generated abstract, and LLM-generated abstract.
@@
Interpretation of the human-generated results of long document summarization. We collected human feedback for generating abstracts of the given scientific papers, denoted as human-generated results. Our participators are well-educated experts in the biology domain, and thus their feedback and interpretation are very important to understand the current difference between LLM preferences and human preferences in summarizing long documents. Following our description in the Dataset section, we selected 10 different papers from the whole testing dataset, and seven participators were numbered H1-H7 in this section. According to Figure 3 (a), the fine-tuned LLM, ChatGLM_lora, can still outperform the best human participator H4 significantly, supported by the p-value computed based on two-sided Wilcoxon rank-sum test (p-value=0.015). Moreover, we demonstrated the performance comparison stratified by different metrics, shown in Figure 3 (b). According to this figure, ChatGLM_lora outperformed other human participators in all of the metrics, which demonstrates that the improvement of fine-tuned LLM is comprehensive. Moreover, we compared the time cost and monetary cost brought by humans and models. Here the cost of machines is computed based on running inference with the fine-tuned ChatGLM model with a standard NVIDIA H100 GPU (per hour price: 0.37\$) [54]. The cost of human participators is computed based on averaging time across different participators based on their reports, and the per-hour price is estimated based on 100,000\$ as annual salary (per hour price: 48.0\$) [55]. According to Extended Data Figure 3 (a), using LLM can reduce both time cost and monetary cost. Overall, our experimental results show that LLMs can generate results that are more similar to the original abstracts after fine-tuning, but they also suffer from increasing variation.
@@
Furthermore, to compare the differences between abstracts from human participators and FMs in detail, we show an example of one abstract in Figure 3 (c). By comparing the original abstract and the human-generated abstract, we found that human-generated outputs have more unique information, highlighted by the blue fonts. The original abstract provides detailed epidemiological data, exact hazard ratios, and a more formal structure. The human-generated report offers a concise overview, underscores the need for broader research, and specifically mentions the lack of a significant finding regarding certain ART regimens. Meanwhile, they also have similarities, as both paragraphs describe the same nationwide South Korean study that investigated outcomes of COVID-19 in People Living with HIV (PLWH) compared to People Living without HIV (PLWoH). They agree that although the crude mortality rate was higher for PLWH, the adjusted difference in mortality was not statistically significant, while the risk of severe or critical illness was clearly elevated. By comparing the original abstract and the LLM-generated abstract, we found that there are substantial numerical inconsistencies between them regarding total cases, proportions of PLWH, comorbidity rates, and mortality. We suspect that these inconsistencies may be due to the wrong numerical understanding made by LLMs when generating the outputs, which might preclude the readers from reaching the correct conclusion. Nevertheless, both paragraphs can describe a nationwide epidemiological study in South Korea over the same time frame, concluding that PLWH is at a higher risk of poor COVID-19 outcomes (either higher incidence or greater severity) and requires dedicated strategies to reduce morbidity and mortality. Therefore, LLMs have advantages in generating an abstract with a more well-structured format and its conclusion is closer to the original description, and human participators have advantages in summarizing the high-level concepts and providing new ideas for including other aspects in the abstract. We also note that the LLMs have some problems with generating data support, so the user needs to analyze the statistics in the output in more detail when using LLMs and make necessary corrections. In Extended Data Figure 1, we show an example of a generated abstract with more textual information, which contains less inconsistent information and is more reliable.
Finally, we explored the understanding gap between fine-tuned LLMs and humans by compressing their outputs from text space into embeddings space based on the OpenAI embedding model, and the cosine similarity between the embeddings using all the selected articles is shown in Extended Data Figure 4 (a). This figure implies that the context information generated by LLMs and humans are similar (nearly all scores are larger than 0.8), and thus LLMs may use skills similar to those of humans to understand and summarize articles so that a person can regard LLM as an assistant when summarizing an article using it.
@@
Figure 4: Evaluations of LMM’s ability to understand scientific figures. (a) Pipeline of our evaluation framework for this task. (b) Performances of different LMMs by the categories of questions. (c) Average performances of different questions by datasets for all selected models. Scores lower than 1.0 are annotated.
Without test-time computation scaling, current LMMs fail to fully understand scientific figures in research papers. In this section, we evaluate the ability of LMMs to understand scientific figures. As humans, to read a scientific figure, we need to understand the overall information of the figure, as well as the detailed information of each panel. Moreover, we may also gain insights or novel biological discoveries. Therefore, we designed questions covering these aspects and selected both open-source (Qwen-VL-Chat [56], LLaVA-13B [57]) and closed-source LMMs (Gemini 1.5 (Flash) [44], GPT-4V, GPT-4o, Claude 3.5 (Sonnet) [45], and o1 [58]) in the evaluation experiment. o1 is known to utilize test-time computation [59] to enhance the reasoning ability, which means that it includes searching strategies when generating the final answer by selecting the correct reasoning path. To compare the outputs of models with ground truth, we computed the accuracy to test whether an LMM can give factual answers without hallucination for all questions. The details of the models and metrics are summarized in the Methods section, and the overview of our evaluation pipeline for this task is illustrated in Figure 4 (a).
According to Figure 4 (b), we found the diversity of the outputs from LMMs across different types of questions based on the boxplot in each panel. o1 ranked as the best model across all types of questions, and open-source models performed obviously worse than the closed-source models in this task. Meanwhile, we found that all methods showed large variance under the questions belonging to insight discovery and color identification. These results show that current LMMs are still struggling in inferring new biology discoveries as well as identifying the color difference. Furthermore, in Figure 4 (c), we visualize the average scores across each dataset for different methods. o1 showed a good performance in answering questions in different categories, highlighting the importance of test-time computation for handling visual reasoning tasks. In detail, we found that o1 still had a large variance in performing tasks related to insight discovery, rank identification, and counting. Furthermore, GPT-4o performed well in analyzing datasets related to RNA-seq (RNA and ZNF71), while was not good at understanding the figure related to method benchmarking analysis (BENCHMARK). Furthermore, we found that GPT-4V showed a strong ability to understand the principle of model design (SPATIALGLUE). Overall, we still need a better model with a stronger ability to understand scientific figures from various sources.
Finally, inspired by [60], we tested if one LMM can generate a reasonable estimate of confidence level. This comparison is important for us to understand the logic of content generation made by these LMMs. Therefore, we requested o1, GPT-4V, and Claude 3.5 (Sonnet) to generate the confidence scores of their answers to all questions, ranging from 0 to 1 and higher values mean better confidence. However, from the scatter plot based on the evaluation score and the confidence score (shown in Extended Data Figure 5), we found that both models did not generate reasonable confidence scores, as the confidence scores do not have statistically significant Spearman correlation with the metric scores (p-value=0.11 for GPT-4V, and p-value=0.14 for Claude 3.5 (Sonnet)). Only the confidence from o1 showed a significant positive correlation with the evaluation score (p-value=0.006), but it still assigns apparently different confidence levels for answers with the same score. Therefore, we also raise concerns about whether to trust the generation of model outputs, and we need a more careful reviewing process for using LMMs as assistants of research figure profiling.

### v2

@@ Abstract
Large language models (LLMs) and large multimodal models (LMMs) promise to accelerate biomedical discovery, yet their reliability remains unclear. We introduce ARIEL (AI Research Assistant for Expert-in-the-Loop Learning), an open-source evaluation and optimization framework that pairs a curated multimodal biomedical corpus with expert-vetted tasks to probe two capabilities: full-length article summarization and fine-grained figure interpretation. Using uniform protocols and blinded PhD-level evaluation, we find that state-of-the-art models generate fluent but incomplete summaries, whereas LMMs struggle with detailed visual reasoning. We later observe that prompt engineering and lightweight fine-tuning substantially improve textual coverage, and a compute-scaled inference strategy enhances visual question answering. We build an ARIEL agent that integrates textual and visual cues, and we show it can propose testable mechanistic hypotheses. ARIEL delineates current strengths and limitations of foundation models, and provides a reproducible platform for advancing trustworthy AI in biomedicine.
@@ 1 Introduction
The success of Foundation Models (FMs) (1), including Large Language Models (LLMs) and Large Multi-Modal Models (LMMs), has been demonstrated across various applications, garnering widespread public interest in research focused on these models (2, 3, 4, 5). A very important target of an FM is to understand various modalities and model them to enable multimodal reasoning (6, 7) as humans. These human skills give us the ability to understand what is being conveyed in a multimodal context, while at the same time being able to externalize it with our linguistic system for processing and refining the content of the text (8). In this work, we investigate how we can utilize FMs to help humans summarize long research papers as well as understand multimodal scientific context.
@@
There are various long documents in the biomedical domain, including examination reports, patient case data, research manuscripts, and others (9). These documents may also include figures as extra information to help us understand the corresponding content. However, due to the explosion of biomedical data nowadays, physicians and researchers are struggling more and more to understand and process such documents in biomedical areas (10). Also, the cost of hiring additional human resources to help process the data is high and therefore not optimal for the biomedical field (11, 12). At the same time, there has been evidence (13, 14, 15) to show that humans tend to engage in complicated reasoning processes and expand their knowledge when answering realistic questions. For example, it is common to use multiple examination results to determine the condition (16) during the diagnosis from a physician. Therefore, how to measure and optimize the reasoning capacity of FMs for complex multimodal data is an important and practical research problem (17). If an FM shows an excellent ability to summarize long texts as well as complicated figures with different panels, it will become a powerful research assistant in this field, which can shorten the time for physicians to make correct diagnoses and for researchers to understand the work of their peers. Previous research has investigated the ability of LLMs to summarize non-medical documents including news (18), codes (19), and reviewer comments (20, 21). Research has also been conducted to evaluate LLMs for summarizing short medical documents including clinical texts (12) and medical evidence (22). Meanwhile, the ability of LMMs to summarize scientific figures has also been explored (23, 24). These papers argue that some advanced models such as GPT-4 can outperform human experts in highlighting the abstract of certain documents or figures, while all models have their limitations due to either hallucinations or out-of-domain queries. Furthermore, most of the recent LLM or LMM evaluations have focused on comparisons among models but ignored the role of humans in evaluations. However, humans are very important references in evaluating the utilities of FMs for their purpose, and understanding the differences between FMs and humans can further ensure the reliability of these FMs for scientific research.
@@
As far as we know, comprehensive human-aware evaluations of LMMs for understanding long biomedical documents and biomedical figures remain unexplored. Although there are several LMM benchmarks on low-level knowledge quizzes, they cannot be applied to our problem due to the complexity of biomedical resources (25, 26, 27). In addition, due to the rapid development of LMMs, we expect more up-to-date evaluations and benchmarks in this time-sensitive domain (28). Most of the existing baselines also do not include human participators as an important metric. Researchers need to get fair and unbiased results carefully to avoid leakage of the test set. Furthermore, inspired by previous work focusing on generating scientific hypotheses from text-level information (29, 30), we are interested in understanding how multimodal biomedical data can help generate new research questions to accelerate scientific progress. Therefore, our LMM evaluations and improvement are challenging and meaningful, which can fill a gap in this advanced area.
@@
Here we present a novel dataset accompanied with analytical framework, named as ARIEL, for evaluating and improving the understanding ability of LLMs and LMMs in the biomedical domain. Here, we focus on two basic tasks, i.e., long manuscript summarization (mainly related to COVID-19 and healthcare) and complicated figure comprehension (mainly related to multiomics) and provide several strategies for performance enhancement. To make a fair and comprehensive comparison, we manually constructed an open-source biomedical article summarization dataset as well as a biomedical figure understanding dataset. Our benchmark based on ARIEL not only systematically evaluates the performance of representative open- and closed-source LLMs and LMMs on both two tasks, but also offers human-aware evaluations from experts in the biomedical field. By analyzing the outputs and experts’ answers, we summarize the strengths and weaknesses of each model type, which enables us to explore how to leverage FMs to facilitate scientific inquires. We also test the performances of ARIEL in generating biomedical hypothesis with different frameworks. Overall, we believe that our multi-functional framework can significantly advance the applications of LLMs and LMMs in the healthcare domain.
@@ 2 Results
Overview of ARIEL To systemically assess the abilities of FMs in understanding research papers, we constructed two datasets: one from manuscripts and another from figures. Data were extracted from PubMed (31) and recently published top-tier journals, with no overlap to the cut-off dates of model training sets. We then developed a general pipeline for generating and automatically evaluating outputs from both closed- and open-source FMs on these datasets (Methods). For text summarization, a language model receives a manuscript and a prompt as input; for figure understanding, a multimodal model takes the figure, its description, and a prompt. In addition to automated metrics, we conducted human evaluation (Human-Eval) with 12 doctoral-level biomedical researchers who judged model outputs and their sources. We then analyzed all experimental results to identify the gaps between current FM capabilities and the requirements of an effective AI research assistant. Finally, we propose initial strategies to leverage FMs for summarizing documents, interpreting figures, and generating hypotheses. A overview of the ARIEL framework is shown in Figure 1.
@@
Figure 1: Landscape of ARIEL. (a) We designed a framework supporting 1. the evaluation of both open-source and closed-source LLMs and LMMs; and 2. the utilization of LLMs and LMMs for helping human researchers resolve both the text summarization task and the figure understanding task with new datasets. We also introduced numerical evaluation metrics and human-aware evaluation metrics for both tasks. Our framework also contains solutions for improving model performances, which are listed in the panel. (b) Number of papers ($n=2571$) with top 10 topics in datasets used for biomedical document summarization based on number ranking. (c) Proportion of questions ($n=88$) in different categories for scientific figure understanding.
Datasets. All questions in our datasets are open-ended. For document summarization, we collected 2571 manuscripts via the PubMed API. Compared with existing benchmarks (32, 33, 34), ARIEL is both both domain-specific and more rigorous, reflecting the complexity of scientific language. We posed open-ended prompts rather than standard QA to increase difficulty. Manuscripts were split into training and test sets, with validation sets derived automatically (35). Topic analysis (Figure 1 (b)) shows dominance of COVID-19 and SARS-CoV-2, followed by healthcare and treatment, reflecting a focus on respiratory epidemics.
For figure understanding, we curated 10 high-quality figures from peer-reviewed publications (24, 25), paired with 88 open-ended questions spanning summarization, color identification, model design, rank identification, recommendation, and insight discovery (Figure 1 (c)). Compared with prior benchmarks (24, 25), our dataset prioritizes low noise and questions requiring reasoning. Human evaluation (Human-Eval) included all figures and a random sample of 10 manuscripts.
@@
Figure 2: Evaluation pipelines and results of text summarization task. We report the average and standard error in all bar plots. (a) Pipeline of evaluation framework for text summarization. (b) Pipeline of evaluation framework for figure understanding (c) Bar plots with standard error of model performances across different evaluation metrics. (d) Bubble plot of averaged model performances (bubble color) and ranks (bubble size) by different evaluation metrics. Models with averaged scores lower than 1.0 are annotated with *.
@@
Large-scale evaluations of long document summarization identify fine-tuned models as better performers. We benchmarked both open-source LLMs (ChatGLM4 (9B) (36), Gemma1 (8B) (37), Llama 3.1 (8B) (38), Mistral (7B) (39), Deepseek-v3 (40), GPT-OSS (20B) (41), GPT-OSS (120B) (41)) and closed-source LLMs (GPT-3.5 (42), GPT-4 (43), GPT-4o (44), Gemini-Flash-1.5 (45), and Claude-Sonnet-3.5 (46)). Performance was assessed with seven established metrics (BLEU, ROUGE-1/2/L, BERTScore, F1-Radgraph, MEDCON)(12, 47, 48, 49, 50, 51, 47). All of the metrics range from 0 to 100 and a higher value means a better output (Figure 2 (a), Methods). Figure 2 (b) shows the pipeline for LMM evaluation, which will be discussed later. To test whether task-specific adaptation improves performance, we fine-tuned the best-performing open-source model, ChatGLM4, using Low-Rank Adaptation (LoRA (52)).
Based on Figure 2 (c), we analyzed the distribution and variance of scores across all evaluation metrics. Overall, open-source models operating in zero-shot inference mode exhibited lower performance than closed-source models. Even advanced open-source systems, such as DeepSeek-V3 and the GPT-OSS series, achieved comparatively modest scores, while models like Gemma1, Llama 3.1, and Mistral often failed to generate meaningful outputs. These shortcomings are likely attributable to data constraints and limited context length. By contrast, ChatGLM4, Claude-Sonnet-3.5, and GPT-4 consistently ranked among the top three models for long-document summarization, aligning with prior studies on the use of LLMs as reviewers (53). Notably, ChatGLM4 surpassed other open-source models across all metrics and demonstrated competitive performance relative to closed-source systems. To further explore its capacity, we applied different prompting strategies and fine-tuned the model (ChatGLM_lora). As shown in Extended Data Figure 2, prompt engineering methods including meta prompting (54) and chain-of-thought reasoning (CoT) (6) had minimal impact on performance. However, fine-tuning substantially improved outcomes, with the adapted model outperforming all closed-source LLMs. This result suggests that domain-specific training can significantly enhance summarization quality and that fine-tuned, smaller-scale models can achieve strong task-specific performance. In Figure 2 (d), we display the average score of each metric for all the models, and the results of both performances and ranks are consistent with conclusions from Figure 2 (c).
@@
Figure 3: Comparisons between human participators and LLMs for the text summarization task. We report the average and standard error in all bar plots. (a) Bar plots of averaged scores and standard errors across all metrics by different participators and ChatGLM_lora. (b) Bar plots of participators’ performances by different evaluation metrics.
@@
Interpretation of the human-generated results of long document summarization. We collected expert feedback on abstract generation for selected scientific papers, denoted as human-generated results. The participants were highly trained specialists in biology, and their evaluations provide critical insights into the divergence between human and LLM preferences when summarizing long documents. As described in the Dataset section, ten papers were sampled from the testing set, and seven participants (denoted H1–H7) contributed feedback. According to Figure 3 (a), the fine-tuned LLM, ChatGLM_lora, can still outperform the best human participator H4 significantly, supported by the p-value computed based on two-sided Wilcoxon rank-sum test (p-value=0.015). Moreover, we demonstrated the performance comparison stratified by different metrics, shown in Figure 3 (b). According to this figure, ChatGLM_lora outperformed other human participators in all of the metrics, which demonstrates that the improvement of fine-tuned LLM is comprehensive. Moreover, we compared the time cost and monetary cost brought by humans and models. Here the cost of machines is computed based on running inference with the fine-tuned ChatGLM model with a standard NVIDIA H100 GPU (per hour price: 0.37\$) (55). The cost of human participators is computed based on averaging time across different participators based on their reports, and the per-hour price is estimated based on 100,000\$ as annual salary (per hour price: 48.0\$) (56). According to Extended Data Figure 3 (a), using LLM can reduce both time cost and monetary cost. Overall, these results indicate that fine-tuned LLMs can generate abstracts more closely aligned with original texts than human experts, but fine-tuned LLMs also show an increase in output variability.
@@
Furthermore, to compare the differences between abstracts from human participators and FMs in detail, we show an example of one abstract in Extended Data Figure 4. By comparing the original abstract and the human-generated abstract, we found that human-generated outputs have more unique information, highlighted by the blue fonts. The original abstract provides detailed epidemiological data, exact hazard ratios, and a more formal structure. The human-generated report offers a concise overview, underscores the need for broader research, and specifically mentions the lack of a significant finding regarding certain ART regimens. Meanwhile, they also have similarities, as both paragraphs describe the same nationwide South Korean study that investigated outcomes of COVID-19 in People Living with HIV (PLWH) compared to People Living without HIV (PLWoH). They agree that although the crude mortality rate was higher for PLWH, the adjusted difference in mortality was not statistically significant, while the risk of severe or critical illness was clearly elevated. By comparing the original abstract and the LLM-generated abstract, we found that there are substantial numerical inconsistencies between them regarding total cases, proportions of PLWH, comorbidity rates, and mortality. We suspect that these inconsistencies may be due to the wrong numerical understanding made by LLMs when generating the outputs, which might preclude the readers from reaching the correct conclusion. Nevertheless, both paragraphs can describe a nationwide epidemiological study in South Korea over the same time frame, concluding that PLWH is at a higher risk of poor COVID-19 outcomes (either higher incidence or greater severity) and requires dedicated strategies to reduce morbidity and mortality. Therefore, LLMs have advantages in generating an abstract with a more well-structured format and its conclusion is closer to the original description, and human participators have advantages in summarizing the high-level concepts and providing new ideas for including other aspects in the abstract. We also note that the LLMs have some problems with generating data support, so the user needs to analyze the statistics in the output in more detail when using LLMs and make necessary corrections. In Extended Data Figure 1, we show an example of a generated abstract with more textual information, which contains less inconsistent information and is more reliable.
Finally, we examined the divergence between fine-tuned LLMs and human outputs by projecting both into an embedding space using the OpenAI embedding model. Cosine similarity scores across all selected articles are presented in Extended Data Figure 5 (a). The results indicate a high degree of alignment, with nearly all similarity values exceeding 0.8, suggesting that LLM-generated content captures contextual information comparable to that produced by humans. This finding implies that LLMs employ strategies resembling human reasoning when processing and summarizing scientific texts, thereby supporting their use as effective assistants in abstract generation.
@@
Figure 4: Evaluations of LMM’s ability to understand scientific figures. We report the average and standard error in all bar plots. (a) Bar plot of accuracy and standard error from different LMMs’ performances by the categories of questions. (c) Bubble plot of average performances (bubble color) and ranks (bubble size) of different questions by datasets for all selected models.
Without test-time computation scaling, current LMMs fail to fully understand scientific figures in research papers. In this section, we assess the capacity of LMMs to interpret scientific figures. Human interpretation of such figures typically involves understanding both the overall context and the detailed content of individual panels, as well as deriving potential insights or biological discoveries. To evaluate these dimensions, we designed a set of questions targeting different levels of comprehension and included both open-source models (Qwen-VL-Chat (57), LLaVA-13B (58)) and closed-source models (Gemini-Flash-1.5 (45), GPT-4V, GPT-4o, Claude-Sonnet-3.5 (46), and o1 (59)). The o1 model, in particular, leverages test-time computation (60) to improve reasoning by employing search strategies to identify the most accurate reasoning path when generating responses.
According to Figure 4 (a), the outputs of LMMs exhibited substantial variability across different question types, as illustrated by the plots in each panel. Among the evaluated models, o1 consistently achieved the highest performance across all question categories, whereas open-source models performed markedly worse than closed-source counterparts. Notably, all models showed considerable variance on tasks involving insight discovery and color identification, indicating persistent limitations in generating novel biological inferences and distinguishing color differences. Figure 4 (b) presents the average scores across datasets for each method. Here, o1 demonstrated robust performance across categories, underscoring the utility of test-time computation in enhancing visual reasoning. Nonetheless, o1 also displayed high variance in tasks including Insight Discovery, Rank Identification, and Counting. GPT-4o performed strongly on datasets involving RNA-seq analyses (RNA and ZNF71) but struggled with figures focused on benchmarking analyses (Benchmark). By contrast, GPT-4V exhibited notable strength in interpreting model design principles, particularly in the SpatialGLUE dataset. Overall, these findings highlight the need for more advanced models with improved capacity to interpret scientific figures across diverse domains and task types.
Finally, inspired by (61), we investigated whether LMMs are capable of producing reliable self-assessments of confidence. This analysis is critical for understanding the internal logic guiding model-generated content. Specifically, we asked o1, GPT-4V, and Claude-Sonnet-3.5 to assign confidence scores (ranging from 0 to 1, with higher values indicating greater confidence) to their answers across all questions. As shown in Extended Data Figure 6, neither GPT-4V nor Claude-Sonnet-3.5 produced meaningful confidence estimates, as their assigned scores were not significantly correlated with evaluation metrics (Spearman correlation, p = 0.11 and p = 0.14, respectively). In contrast, o1 exhibited a significant positive correlation between confidence and accuracy (p = 0.006), although inconsistencies remained, with the model assigning divergent confidence levels to responses with identical scores. Overall, these findings raise concerns about the reliability of confidence estimation in current LMMs and underscore the need for careful human oversight when employing them as assistants for scientific figure interpretation.